Given any positive even integer, x, the positive difference between the smallest odd number greater than 7x − 2 and the largest odd number less than 3x + 5 can be written in the form ax + b. What is a + b?
Answer by RAY100(1637)

Given x is an even integer,,,7x-2 is also even,,,smallest odd integer larger is 7x-2 +1 = 7x-1
.
3x+5 is odd integer,,,,,largest odd number less is,,,,,3x+5 -2 = 3x+3
.
.
.
(7x-1)-(3x+3) = 4x -4,,,,,ax+b
.
a=4,,,,b=-4,,,,,,,a+b = 0
.